OpenStudy (anonymous):

if f(x)=3x+4 and (x)=x+2, find f-g

OpenStudy (anonymous):

f - g = (3x+4) - (x+2) = 3x + 4 - x - 2 = 2x + 2 Kepp in mind that when you perform -(x+2) you are negating what is inside the parenthesis.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome.